# GEDAN: Learning the Edit Costs for Graph Edit Distance

### v1

@@ Abstract
Graph Edit Distance (GED) is defined as the minimum cost transformation of one graph into another and is a widely adopted metric for measuring the dissimilarity between graphs. The major problem of GED is that its computation is NP-hard, which has in turn led to the development of various approximation methods, including approaches based on neural networks (NN). Most of these NN-based models simplify the problem of GED by assuming unit-cost edit operations—a rather unrealistic constraint in real-world applications. In this work, we present a novel Graph Neural Network framework that approximates GED using both supervised and unsupervised training. In the unsupervised setting, it employs a gradient-only self-organizing mechanism that enables optimization without ground-truth distances. Moreover, a core component of our architecture is the integration of a Generalized Additive Model, which allows the flexible and interpretable learning of context-aware edit costs. Experimental results show that the proposed method achieves similar results as state-of-the-art reference methods, yet significantly improves both adaptability and interpretability. That is, the learned cost function offers insights into complex graph structures, making it particularly valuable in domains such as molecular analysis and structural pattern discovery.
@@ 1 Introduction and Related Work
Graphs are a versatile data structure that can be used to represent both the features of entities and the relationships between them. Due to their power and flexibility, graphs have found numerous applications. For example, graph-based representations are successfully used in the analysis of social networks (Chen, 2013), the recognition of human poses (Mathis et al., 2018), the modeling of road infrastructure (Haklay and Weber, 2008) or the representation of watercourses (Fankhauser et al., 2024). Graphs can also be used in bio-informatics and computational chemistry, for example, to represent proteins (Bernstein et al., 1977) or molecules (Bento et al., 2020), the field we consider as possible application in this paper.
@@
In the last decades, several methods have been proposed and researched to assess the dissimilarity or similarity between graphs, such as diverse graph kernels or graph matching procedures Conte et al. (2004). Among these, Graph Edit Distance (GED) Gao et al. (2010) can be interpreted as a standard approach. Basically, GED measures the distance between two graphs in terms of the minimum cost operations required to transform one graph into the other. The considered operations typically include insertions, deletions and substitutions of both nodes and edges. One of the most important advantages of GED is that it can be applied to virtually all types of graphs and thus, GED can be flexibly adapted to any graph-based application. However, exact computation of GED is an NP-hard problem (Aho et al., 1974), which makes its application to larger graphs, or in scenarios requiring a large number of comparisons, impractical.
@@
To address this limitation and to reduce the computational complexity of GED, several strategies for relaxing the problem have been proposed. These techniques range from binary linear programming (Justice and III, 2006), over fast suboptimal algorithms preserving classification accuracy (Neuhaus et al., 2006; Riesen and Bunke, 2009), to quadratic assignment models Bougleux et al. (2017) and local search algorithms Boria et al. (2020). More recently, diverse approaches to learn and approximate GED using Graph Neural Networks (GNNs) have been proposed (Bai et al., 2019; Ranjan et al., 2022; Zhuo and Tan, 2022). A very recent approach Jain et al. (2024), also based on GNNs, differs from the others in explicitly formulating the GED as a matrix optimization problem. That is, this method approximates the GED by learning the permutation that minimizes the assignment cost between the substructures of two graphs.
Since the GNN-based methods for learning GED are trained in a supervised manner, a ground truth is required for the distances between the training graphs. Clearly, this is a significant limitation as the computation of the actual distances between graphs stemming from large datasets is computationally prohibitive, limiting the practical applicability of these methods. This contradiction was recently pointed out by Yang et al. (2024), who proposed strategies such as transfer learning and few-shot learning to mitigate the problem.
A second problem is that the vast majority of the GNN-based approaches assume unit costs for the individual operations. This can be regarded as particularly restrictive in graph-based applications in chemistry or bio-informatics, where nodes and edges of graphs are often used to represent atoms and chemicals bonds, respectively. In this context, it is known that the substitution of atoms, or functional groups, does not occur at uniform cost, but depends on specific structural and functional factors Guilbert et al. (2021); Edzards et al. (2024).
Our paper addresses the two challenges mentioned above separately. First, based on the idea proposed by Jain et al. (2024), our novel model combines GNNs Wu et al. (2021) with Generalized Additive Models (GAMs) McCaffrey (1992) and approximates the GED by a soft-assignment obtained with a Gumbel-Sinkhorn network Mena et al. (2018). A key advantage of this novel approach is that it can be used in unsupervised contexts, eliminating the need for explicit ground truth distances. Second, unlike previous methods, our approach learns edit costs directly through gradient descent, making the GED consistent with respect to graph properties. The benefit of this approach is that the learned costs might reveal lesser-known relationships between the underlying graphs and provide new insights for domain experts.
@@
The remaining paper is structured as follows. In Section 2, we describe both the method used to approximate the GED and the method to learn the costs of the edit operations. In Section 3, we present the results of our experimental evaluation and present a use case of our novel approach stemming from the field of computational chemistry. In Section 4, we conclude the paper and propose future research work directions.
@@ 2 Methods
In the present paper, we define a graph as a triple $G=(V,E,\mu)$, where $V$ is the set of nodes, $E\subseteq\{(u,v)|u,v\in V,u\neq v\}$ is the set of edges, and $\mu:V\rightarrow L$ is a function that assigns a label $\mu(v)=l\in L$ to each node $v\in V$ (where $L$ is a given label alphabet). Sometimes this definition is complemented with a second function, which also assigns a label to the edges in the graph. In this paper, however, we only use node-labeled graphs (in fact, the methods proposed in this section could also be adapted to labeled edges).
The overall aim of our novel method is to compute the Graph Edit Distance GED($G,G^{\prime}$) between two graphs $G$ and $G^{\prime}$. Our approach is based on a reformulation of the problem of GED to a Quadratic Assignment Problem (QAP) Bougleux et al. (2017) as proposed in Jain et al. (2024). However, the method proposed in this paper (illustrated in Fig. 1) differs significantly from the previous approaches in four key aspects, which are discussed in detail in the following four subsections.
@@
Figure 1: Structure of the GEDAN architecture. Learnable layers are shown in blue color, static or frozen layers are shown in yellow color.
@@ 2.1 Penalization with Message Passing
where $A^{k}$ is an injective neighborhood aggregation function that captures information up to the $k$-hop neighborhood of node $v$, while preserving its structural uniqueness. In our implementation, we instantiate $A^{k}$ as the Graph Isomorphism Network (GIN) Xu et al. (2019), which generalizes the Weisfeiler-Lehman test Shervashidze et al. (2011) and is known for its expressive power in distinguishing graph structures.

### v2

@@ Abstract
Graph Edit Distance (GED) is defined as the minimum cost transformation of one graph into another and is a widely adopted metric for measuring the dissimilarity between graphs. The major problem of GED is that its computation is NP-hard, which has in turn led to the development of various approximation methods, including approaches based on neural networks (NN). However, most NN methods assume a unit cost for edit operations – a restrictive and often unrealistic simplification, since topological and functional distances rarely coincide in real-world data. In this paper, we propose a fully end-to-end Graph Neural Network framework for learning the edit costs for GED, at a fine-grained level, aligning topological and task-specific similarity. Our method combines an unsupervised self-organizing mechanism for GED approximation with a Generalized Additive Model that flexibly learns contextualized edit costs. Experiments demonstrate that our approach overcomes the limitations of non–end-to-end methods, yielding directly interpretable graph matchings, uncovering meaningful structures in complex graphs, and showing strong applicability to domains such as molecular analysis.
@@ 1 Introduction and Related Work
Graphs are a versatile data structure that can be used to represent both the features of entities and the relationships between them. Due to their power and flexibility, graphs have found numerous applications. For example, graph-based representations are successfully used in the analysis of social networks (Chen, 2013), the recognition of human poses (Mathis et al., 2018) or the representation of water bodies (Fankhauser et al., 2024). Graphs can also be used in bio-informatics and computational chemistry, for example, to represent proteins (Bernstein et al., 1977) or molecules (Bento et al., 2020), the field we consider as possible application in this paper.
@@
In the last decades, several methods have been proposed and researched to assess the dissimilarity or similarity between graphs, such as diverse graph kernels or graph matching procedures (Conte et al., 2004; Vento, 2013). Among these, Graph Edit Distance (GED) (Bunke & Allermann, 1983) can be interpreted as a standard approach. Basically, GED measures the distance between two graphs in terms of the minimum cost operations required to transform one graph into the other. The considered operations typically include insertions, deletions, and substitutions of both nodes and edges. One of the most important advantages of GED is that it can be applied to virtually all types of graphs and thus, GED can be flexibly adapted to any graph-based application. However, exact computation of GED is an NP-hard problem (Bunke, 1997), which makes its application to larger graphs, or in scenarios requiring a large number of comparisons, impractical.
@@
To address this limitation and to reduce the computational complexity of GED, several strategies for relaxing the problem have been proposed. These techniques range from binary linear programming (Justice & III, 2006), over fast suboptimal algorithms preserving classification accuracy (Neuhaus et al., 2006b; Riesen & Bunke, 2009b), to quadratic assignment models (Bougleux et al., 2017) and local search algorithms (Boria et al., 2020). More recently, diverse approaches to learn and approximate GED using Graph Neural Networks (GNNs) have been proposed (Bai et al., 2019a; Ranjan et al., 2022; Zhuo & Tan, 2022). A very recent approach (Jain et al., 2024), also based on GNNs, differs from the others in explicitly formulating the GED as a matrix optimization problem. That is, this method approximates the GED by learning the permutation that minimizes the assignment cost between the substructures of two graphs.
Most existing work approaches GED primarily as an approximation problem for measuring graph dissimilarity, typically assuming fixed – often uniform – edit costs. This simplification is unrealistic, since treating all edits as equally costly ignores the fact that structural similarity is not necessarily aligned with functional or task-specific similarity. In molecular graphs, for example, the substitution of a single atom (e.g., replacing a hydrogen with a chlorine) can drastically alter pharmacological activity, toxicity, or binding affinity. Conversely, two molecules that are topologically quite dissimilar may nonetheless exhibit comparable biological properties – a phenomenon known as scaffold hopping in drug design (Sun et al., 2012).
Consequently, adapting GED to reflect meaningful, task-specific edit costs is non-trivial, since most available methods are supervised and rely on pre-computed distances. To make GED reflect functional similarity, it is necessary to learn the underlying edit costs – quantities that are not known a priori but constitute the very target of learning. While unsupervised methods such as GraphCL (You et al., 2020) or UGraphEmb (Bai et al., 2019b) can learn structural representations through augmentation, they cannot explicitly capture the actual edit costs. Recent methods such as EPIC (Heo et al., 2024) and GMSM (Pellizzoni et al., 2024) attempt to learn edit costs. EPIC employs them primarily for data augmentation, while GMSM restricts node-specific costs to the final matching step, thereby missing multi-scale granularity.
In this paper, we address the problem of directly learning edit costs that enable GED to align with task-relevant properties. Building on the approach of Jain et al. (2024), we combine GNNs with Generalized Additive Models (GAM) (McCaffrey, 1992), and employ a Gumbel-Sinkhorn network (Mena et al., 2018) for soft GED approximation. This approximation is performed in an unsupervised, self-organizing manner driving the model to restructure itself to minimize the overall assignment cost. Finally, we apply contrastive learning to align structural and functional distances between graphs, thereby improving this alignment while providing interpretable insights into graph relationships.
@@
The remaining paper is structured as follows. In Section 2, we describe both the method used to approximate the GED and the method to learn the costs of the edit operations. In Section 3, we present the results of our experimental evaluation and present a use case of our novel approach stemming from the field of computational chemistry. In Section 4, we conclude the paper by discussing its limitations and outlining directions for future research.
@@ 2 Methods
In the present paper, we define a graph as a triple $G=(V,E,\mu)$, where $V$ is the set of nodes, $E\subseteq\{(u,v)|u,v\in V,u\neq v\}$ is the set of edges, and $\mu:V\rightarrow L$ is a function that assigns a label $\mu(v)=l\in L$ to each node $v\in V$ (where $L$ is a given label alphabet). Hence, we focus on node-labeled graphs, though our approach can be extended to include edge labels.
Our goal is not merely to compute the GED, but to learn a set of edit costs $\mathcal{C}$ such that GED aligns with a task-specific functional distance $d_{f}$. Formally, given three graphs $G,G^{\prime},G^{\prime\prime}$, we aim to enforce
$d_{f}(G,G^{\prime})<d_{f}(G,G^{\prime\prime})\implies GED(G,G^{\prime};\mathcal{C})<GED(G,G^{\prime\prime};\mathcal{C}),$ (1)
ensuring that as functional distance increases, so does the corresponding GED. Consequently, the set of edit costs $\mathcal{C}$ is not fixed, but learned to preserve the ordering induced by $d_{f}$.
To approximate GED, our approach is based on a reformulation of the problem of GED to a Quadratic Assignment Problem (QAP) (Bougleux et al., 2017) as proposed in Jain et al. (2024). However, the method proposed in this paper differs from prior work in four key aspects, which we detail in the following subsections.
@@ 2.1 Penalization with Message Passing
where $A^{k}$ is an injective neighborhood aggregation function that captures information up to the $k$-hop neighborhood of node $v$, while preserving its structural uniqueness. In our implementation, we instantiate $A^{k}$ as the Graph Isomorphism Network (GIN) (Xu et al., 2019), which generalizes the Weisfeiler-Lehman test (Shervashidze et al., 2011) and is known for its expressive power in distinguishing graph structures.